3 .What amount should be report on the current year’s income statement for insurance expense?

A+T Williamson Company is making adjustment entries for the year ended December 31 of the current year. In developing information for the adjusting entries, the accountant learned the following.


a. A two-year insurance premium of $4800 was paid on October 1 of the current year for coverage beginning on that date. The bookkeeper debited the full amount to Prepaid Insurance in October 1.

b. At December 31 of the current year, the following data relating to Shipping Supplies were obtained from the records and supporting documents.


Shipping supplies on hand, January 1 of the year current year                          $13,000

Purchases of shipping supplies during the current year                                      $75,000

Skipping supplies on hand, counted on December 31 of the current year      $ 20,000



Required

3 . What amount should be report on the current year’s income statement for insurance expense?

Do not round intermediate calculation

Insurance expense                  { ? }

Shipping Supplies expense    { ? }
Answer #1

Calculate amount

a) Insurance expense = 4800/24*3 = $600

b) Shipping supplies expense = 13000+75000-20000 = $68000